# Second FRCSyn-onGoing: Winning Solutions and Post-Challenge Analysis to Improve Face Recognition with Synthetic Data

### v1

@@ IV Second FRCSyn-onGoing: Systems Description
Figure 10: Framework proposed by the CBSR-Samsung team.Figure 11: Framework proposed by the BOVIFOCR-UFPR team.
TABLE IV: Results of the teams that ranked among the top-6 in at least one sub-task, ordered by the average rank in all the sub-tasks. For each team, we report the ranking metric and the position across all the sub-tasks. The best result of each sub-task is highlighted in bold. We mark with a ‘-’ if the team did not participate in a sub-task. TO = Trade-off, AVG = average accuracy
Task 1: Bias Mitigation
Task 2: Overall Improvement
Team
Task 1.1
TO [%]
Task 1.2
TO [%]
Task 1.3
TO [%]
Task 2.1
AVG [%]
Task 2.2
AVG [%]
Task 2.3
AVG [%]
ADMIS
94.30 (2)
95.72 (2)
96.50 (1)
91.19 (3)
92.92 (2)
94.15 (5)
OPDAI
93.75 (4)
94.12 (3)
95.96 (4)
91.93 (1)
92.04 (3)
95.23 (2)
ID R&D
96.73 (1)
96.73 (1)
86.73 (8)
91.86 (2)
91.86 (4)
94.05 (6)
K-IBS-DS
92.91 (6)
93.72 (5)
96.17 (2)
91.05 (4)
91.61 (5)
95.42 (1)
CTAI
93.21 (5)
93.21 (6)
95.41 (7)
90.59 (5)
90.59 (6)
94.56 (3)
Idiap-SynthDistill
89.70 (9)
93.50 (1)
INESC-IGD
92.28 (7)
94.05 (4)
95.65 (5)
83.16 (8)
85.40 (7)
89.43 (8)
UNICA-IGD-LSI
91.89 (8)
91.89 (7)
96.00 (3)
87.80 (7)
87.80 (8)
92.79 (7)
SRCN_AIVL
94.06 (3)
CBSR-Samsung
95.57 (6)
94.20 (4)
BOVIFOCR-UFPR
90.48 (8)
89.97 (6)
@@ V Second FRCSyn-onGoing: Results
Next, we describe in Sections V-A and V-B the main results achieved in Tasks 1 and 2, respectively. These results are further analyzed in Section V-C focusing on specific demographic groups and individual databases. Finally, we discuss in Section V-D common trends among the different teams and compare the results with those obtained in the 1 edition. We present in Table IV the ranking and key results of the 2 FRCSyn-onGoing.
@@
Table V shows the results achieved by participants in Task 1, focused on demographic bias mitigation. Teams are ranked by descending order of TO, which tends to correlate to the ascending order of SD (i.e., from less to more biased FR systems). Notably, the winner of Sub-Tasks 1.1 and 1.2, ID R&D (96.73% TO), demonstrates a significant negative GAP value (-5.31%), showing a higher performance when training the FR system with synthetic data compared to real data (i.e., CASIA-WebFace [42]). Furthermore, in Sub-Task 1.1, we can observe that teams with negative GAP values considered Diffusion Models for the generation of synthetic data (i.e., ID R&D uses HDT, SRCN_AIVL combines DCFace and IDiff-Face, and CTAI combines DCFace and GANDiffFace), showing that this generation method may work better than real data in scenarios with limited data. Next, after removing the limitation on the number of synthetic images (i.e., Sub-Task 1.2), the TO value of most FR systems increases, which leads to performance and fairness improvement simultaneously. For instance, for the ADMIS team (ranked top-2 in both Sub-Task 1.1 and 1.2), the TO value increases to 95.72% in Sub-Task 1.2 (i.e., 1.42% TO improvement compared to Sub-Task 1.1). Also, the GAP value decreases from 1.47% to -0.56%, obtaining better results when increasing the amount of synthetic data in comparison to limited real data (i.e., CASIA-WebFace). Another example is the OPDAI team, which raised from top-4 to top-3 positions between Sub-Task 1.1 and 1.2. Its TO value increases to 94.12% (i.e., 0.37% TO improvement from Sub-Tasks 1.1 to 1.2), and the GAP value is reduced from 1.02% to 0.71%. These findings emphasize the potential of generating large number of synthetic face images from different demographic groups to mitigate bias in existing FR technology. Finally, we analyze in Sub-Task 1.3 the case of using both, real and synthetic data, in the FR training process. In general, we can observe considerable improvements in terms of TO values, along with higher negative GAP values for all the top-6 teams, e.g., ADMIS (96.50% TO, -1.33 GAP), K-IBS-DS (96.17% TO, -1.37% GAP), and UNICA-IGD-LSI (96.00% TO, -5.33% GAP). These results highlight that the combination of synthetic and real data achieves superior FR performance compared to the case of only training with real data. Finally, it is noteworthy to compare the best results achieved in Sub-Task 1.2, i.e., unconstrained synthetic data, and Sub-Task 1.3, i.e., constrained synthetic + real data. The ID R&D team achieves 96.73% TO in Sub-Task 1.2, whereas ADMIS achieves 96.50% TO in Sub-Task 1.3, showing that unlimited synthetic data for training can even outperform FR systems trained with limited synthetic + real data. These results motivate the use of synthetic data for demographic bias mitigation, improving at the same time privacy as no real identities are seen by the network.
TABLE V: Ranking for the three sub-tasks considered in Task 1. For each sub-task, we highlight in bold the best team according to the Trade-Off. TO = Trade-Off, AVG = Average accuracy, SD = Standard Deviation of accuracy, FNMR = False Non-Match Rate, FMR = False Match Rate, AUC = Area Under Curve, GAP = Gap to Real.
Sub-Task 1.1 (Bias Mitigation): Synthetic Data (Constrained)
Pos.
Team
TO [%]
AVG [%]
SD [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
ID R&D
96.73
97.55
0.82
3.17 (1)
99.57 (1)
-5.31
2
ADMIS
94.30
95.10
0.80
11.38 (3)
98.96 (3)
1.47
3
SRCN_AIVL
94.06
95.12
1.07
10.72 (2)
98.83 (4)
-0.54
4
OPDAI
93.75
94.92
1.17
11.85 (4)
99.51 (2)
1.02
5
CTAI
93.21
94.74
1.53
14.38 (5)
98.33 (6)
-0.63
6
K-IBS-DS
92.91
94.11
1.2
15.03 (6)
98.47 (5)
1.58
Sub-Task 1.2 (Bias Mitigation): Synthetic Data (Unconstrained)
Pos.
Team
TO [%]
AVG [%]
SD [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
ID R&D
96.73
97.55
0.82
3.17 (1)
99.57 (2)
-5.31
2
ADMIS
95.72
96.50
0.78
6.33 (2)
99.51 (3)
-0.56
3
OPDAI
94.12
95.22
1.11
10.78 (3)
98.92 (4)
0.71
4
INESC-IGD
94.05
95.22
1.17
11.03 (4)
98.70 (5)
1.04
5
K-IBS-DS
93.72
94.88
1.16
12.75 (5)
99.66 (1)
0.77
6
CTAI
93.21
94.74
1.53
14.38 (6)
98.33 (6)
-0.63
Sub-Task 1.3 (Bias Mitigation): Synthetic + Real Data (Constrained)
Pos.
Team
TO [%]
AVG [%]
SD [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
ADMIS
96.50
97.25
0.75
3.90 (1)
99.72 (1)
-1.33
2
K-IBS-DS
96.17
96.92
0.75
5.88 (4)
99.54 (2)
-1.37
3
UNICA-IGD-LSI
96.00
96.70
0.70
5.90 (5)
99.49 (3)
-5.33
4
OPDAI
95.96
96.80
0.84
4.90 (2)
99.54 (2)
-0.03
5
INESC-IGD
95.65
96.33
0.67
6.15 (6)
99.18 (5)
-0.12
6
CBSR-Samsung
95.57
96.54
0.97
5.00 (3)
99.41 (4)
-24.43
TABLE VI: Ranking for the three sub-tasks considered in Task 2. For each sub-task, we highlight in bold the best team according to the Average Accuracy. AVG = Average accuracy, FNMR = False Non-Match Rate, FMR = False Match Rate, AUC = Area Under Curve, GAP = Gap to Real.
Sub-Task 2.1 (Overall Improvement): Synthetic Data (Constrained)
Pos.
Team
AVG [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
OPDAI
91.93
17.63 (2)
97.30 (2)
3.09
2
ID R&D
91.86
10.36 (1)
97.48 (1)
2.99
3
ADMIS
91.19
20.41 (3)
97.04 (3)
2.78
4
K-IBS-DS
91.05
24.87 (6)
96.09 (6)
2.60
5
CTAI
90.59
21.88 (4)
96.40 (5)
-1.94
6
BOVIFOCR-UFPR
89.97
24.04 (5)
96.70 (4)
3.71
Sub-Task 2.2 (Overall Improvement): Synthetic Data (Unconstrained)
Pos.
Team
AVG [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
Idiap-SynthDistill
93.50
9.17 (1)
97.17 (4)
-0.05
2
ADMIS
92.92
14.45 (3)
97.76 (1)
0.21
3
OPDAI
92.04
16.48 (4)
97.41 (3)
3.00
4
ID R&D
91.86
10.36 (2)
97.48 (2)
2.99
5
K-IBS-DS
91.61
22.48 (6)
96.54 (5)
1.96
6
CTAI
90.59
21.88 (5)
96.40 (6)
-1.94
Sub-Task 2.3 (Overall Improvement): Synthetic + Real Data (Constrained)
Pos.
Team
AVG [%]
FNMR@FMR=1%
AUC [%]
GAP [%]
1
K-IBS-DS
95.42
9.49 (5)
98.14 (6)
-2.15
2
OPDAI
95.23
7.54 (1)
98.70 (1)
-0.52
3
CTAI
94.56
8.85 (4)
98.41 (3)
-6.01
4
CBSR-Samsung
94.20
8.62 (3)
98.17 (4)
-4.40
5
ADMIS
94.15
10.99 (6)
98.46 (2)
-1.10
6
ID R&D
94.05
8.00 (2)
98.16 (5)
0.07
Figure 12: DET curves of Sub-Tasks 1.1 (left), 1.2 (Middle), and 1.3 (right), including the results achieved for the top-1 team (i.e., ID R&D, ID R&D, ADMIS, respectively) in each demographic group.
@@
Table VI provides the results achieved by participants in Task 2, focusing not only on bias mitigation but also other challenges in FR such as age, pose, and occlusions. Teams are ranked in descending order based on the average verification accuracy across the four databases. Notably, in all sub-tasks, the AVG is lower than the achieved in Task 1 for the BUPT-BalancedFace [43] database, showing the additional challenges introduced by the AgeDB [44], CFP-FP [45], and ROF [46] databases. This trend can also be observed in GAP results, which tend to be worse for Sub-Tasks 2.1 and 2.2 compared to Sub-Tasks 1.1 and 1.2, suggesting that it is far more difficult to emulate the conditions of these real databases with synthetic data by itself. For example, in Sub-Task 2.1, although the top-3 teams achieve high AVG results, they exhibit a considerable positive GAP value (i.e., OPDAI 91.93% AVG, 3.09% GAP; ID R&D 91.86% AVG, 2.99% GAP; ADMIS 91.19% AVG, 2.78% GAP), showing that a FR model trained only with real data (i.e., CASIA-WebFace) adapts better to adverse image conditions such as aging, pose, or occlusions. Focusing on Sub-Task 2.2, the team ranked top-1, i.e., Idiap-SynthDistill, achieves much better results compared to the best result of Sub-Task 2.1 (i.e., 93.50% vs. 91.93% AVG), proving that unlimited synthetic data can further improve the performance of the system. Finally, in Sub-Task 2.3, most teams report better AVG and higher negative GAP values (e.g., K-IBS-DS achieves 95.42% AVG, -2.15% GAP), proving again that synthetic data combined with real data can alleviate existing limitations within FR technology.

### v2

@@ IV Second FRCSyn-onGoing: Systems Description
Task 1: Bias Mitigation | Task 2: Overall Improvement | |||||||||||||||||
Team |
ADMIS | 94.30 (2) | 95.72 (2) | 96.50 (1) | 91.19 (3) | 92.92 (2) | 94.15 (5) | ||||||||||||
OPDAI | 93.75 (4) | 94.12 (3) | 95.96 (4) | 91.93 (1) | 92.04 (3) | 95.23 (2) | ||||||||||||
ID R&D | 96.73 (1) | 96.73 (1) | 86.73 (8) | 91.86 (2) | 91.86 (4) | 94.05 (6) | ||||||||||||
K-IBS-DS | 92.91 (6) | 93.72 (5) | 96.17 (2) | 91.05 (4) | 91.61 (5) | 95.42 (1) | ||||||||||||
CTAI | 93.21 (5) | 93.21 (6) | 95.41 (7) | 90.59 (5) | 90.59 (6) | 94.56 (3) | ||||||||||||
Idiap-SynthDistill | - | 89.70 (9) | - | - | 93.50 (1) | - | ||||||||||||
INESC-IGD | 92.28 (7) | 94.05 (4) | 95.65 (5) | 83.16 (8) | 85.40 (7) | 89.43 (8) | ||||||||||||
UNICA-IGD-LSI | 91.89 (8) | 91.89 (7) | 96.00 (3) | 87.80 (7) | 87.80 (8) | 92.79 (7) | ||||||||||||
SRCN_AIVL | 94.06 (3) | - | - | - | - | - | ||||||||||||
CBSR-Samsung | - | - | 95.57 (6) | - | - | 94.20 (4) | ||||||||||||
BOVIFOCR-UFPR | - | 90.48 (8) | - | 89.97 (6) | - | - |
Sub-Task 1.1 (Bias Mitigation): Synthetic Data (Constrained) | |||||||
Pos. | Team | TO [%] | AVG [%] | SD [%] | FNMR@FMR=1% | AUC [%] | GAP [%] |
1 | ID R&D | 96.73 | 97.55 | 0.82 | 3.17 (1) | 99.57 (1) | -5.31 |
2 | ADMIS | 94.30 | 95.10 | 0.80 | 11.38 (3) | 98.96 (3) | 1.47 |
3 | SRCN_AIVL | 94.06 | 95.12 | 1.07 | 10.72 (2) | 98.83 (4) | -0.54 |
4 | OPDAI | 93.75 | 94.92 | 1.17 | 11.85 (4) | 99.51 (2) | 1.02 |
5 | CTAI | 93.21 | 94.74 | 1.53 | 14.38 (5) | 98.33 (6) | -0.63 |
6 | K-IBS-DS | 92.91 | 94.11 | 1.2 | 15.03 (6) | 98.47 (5) | 1.58 |
Sub-Task 1.2 (Bias Mitigation): Synthetic Data (Unconstrained) | |||||||
Pos. | Team | TO [%] | AVG [%] | SD [%] | FNMR@FMR=1% | AUC [%] | GAP [%] |
1 | ID R&D | 96.73 | 97.55 | 0.82 | 3.17 (1) | 99.57 (2) | -5.31 |
2 | ADMIS | 95.72 | 96.50 | 0.78 | 6.33 (2) | 99.51 (3) | -0.56 |
3 | OPDAI | 94.12 | 95.22 | 1.11 | 10.78 (3) | 98.92 (4) | 0.71 |
4 | INESC-IGD | 94.05 | 95.22 | 1.17 | 11.03 (4) | 98.70 (5) | 1.04 |
5 | K-IBS-DS | 93.72 | 94.88 | 1.16 | 12.75 (5) | 99.66 (1) | 0.77 |
6 | CTAI | 93.21 | 94.74 | 1.53 | 14.38 (6) | 98.33 (6) | -0.63 |
Sub-Task 1.3 (Bias Mitigation): Synthetic + Real Data (Constrained) | |||||||
Pos. | Team | TO [%] | AVG [%] | SD [%] | FNMR@FMR=1% | AUC [%] | GAP [%] |
1 | ADMIS | 96.50 | 97.25 | 0.75 | 3.90 (1) | 99.72 (1) | -1.33 |
2 | K-IBS-DS | 96.17 | 96.92 | 0.75 | 5.88 (4) | 99.54 (2) | -1.37 |
3 | UNICA-IGD-LSI | 96.00 | 96.70 | 0.70 | 5.90 (5) | 99.49 (3) | -5.33 |
4 | OPDAI | 95.96 | 96.80 | 0.84 | 4.90 (2) | 99.54 (2) | -0.03 |
5 | INESC-IGD | 95.65 | 96.33 | 0.67 | 6.15 (6) | 99.18 (5) | -0.12 |
6 | CBSR-Samsung | 95.57 | 96.54 | 0.97 | 5.00 (3) | 99.41 (4) | -24.43 |
@@ V Second FRCSyn-onGoing: Results
Next, we describe in Sections V-A and V-B the main results achieved in Tasks 1 and 2, respectively. These results are further analyzed in Section V-C focusing on specific demographic groups and individual databases. Finally, we discuss in Section V-D common trends among the different teams and compare the results with those obtained in the 1 edition. We present in Table V the ranking and key results of the 2 FRCSyn-onGoing.
Team | Model |
1.1 1.2 | ID R&D | Proposed | 96.80 | 96.90 | 97.20 | 97.80 | 98.10 | 99.20 | 96.50 | 97.90 | 97.55 | 0.82 | 96.73 | -5.31 | ||||||||||||||||||||||||||
Baseline | 85.70 | 90.00 | 94.10 | 95.20 | 87.10 | 90.80 | 97.60 | 98.40 | 92.36 | 4.40 | 87.96 | |||||||||||||||||||||||||||||
1.3 | ADMIS | Proposed | 96.20 | 97.30 | 97.30 | 97.10 | 96.20 | 97.40 | 98.60 | 97.90 | 97.25 | 0.75 | 96.50 | -1.33 | ||||||||||||||||||||||||||
Baseline | 92.90 | 94.60 | 97.00 | 98.00 | 93.70 | 96.90 | 97.10 | 97.50 | 95.96 | 1.81 | 94.15 |
@@ V-A Task 1: Bias Mitigation
Table VI shows the results achieved by participants in Task 1, focused on demographic bias mitigation. Teams are ranked by descending order of TO, which tends to correlate to the ascending order of SD (i.e., from less to more biased FR systems). Notably, the winner of Sub-Tasks 1.1 and 1.2, ID R&D (96.73% TO), demonstrates a significant negative GAP value (-5.31%), showing a higher performance when training the FR system with synthetic data compared to real data (i.e., CASIA-WebFace [51]). Furthermore, in Sub-Task 1.1, we can observe that teams with negative GAP values considered Diffusion Models for the generation of synthetic data (i.e., ID R&D uses HDT, SRCN_AIVL combines DCFace and IDiff-Face, and CTAI combines DCFace and GANDiffFace), showing that this generation method may work better than real data in scenarios with limited data. Next, after removing the limitation on the number of synthetic images (i.e., Sub-Task 1.2), the TO value of most FR systems increases, which leads to performance and fairness improvement simultaneously. For instance, for the ADMIS team (ranked top-2 in both Sub-Task 1.1 and 1.2), the TO value increases to 95.72% in Sub-Task 1.2 (i.e., 1.42% TO improvement compared to Sub-Task 1.1). Also, the GAP value decreases from 1.47% to -0.56%, obtaining better results when increasing the amount of synthetic data in comparison to limited real data (i.e., CASIA-WebFace). Another example is the OPDAI team, which raised from top-4 to top-3 positions between Sub-Task 1.1 and 1.2. Its TO value increases to 94.12% (i.e., 0.37% TO improvement from Sub-Tasks 1.1 to 1.2), and the GAP value is reduced from 1.02% to 0.71%. These findings emphasize the potential of generating large number of synthetic face images from different demographic groups to mitigate bias in existing FR technology. Finally, we analyze in Sub-Task 1.3 the case of using both, real and synthetic data, in the FR training process. In general, we can observe considerable improvements in terms of TO values, along with higher negative GAP values for all the top-6 teams, e.g., ADMIS (96.50% TO, -1.33 GAP), K-IBS-DS (96.17% TO, -1.37% GAP), and UNICA-IGD-LSI (96.00% TO, -5.33% GAP). Moreover, Table VII shows the performance of each demographic group for the baseline models, i.e., those trained only with real data, and the models proposed by the top-ranked teams for each of the sub-tasks of Task 1. The ID R&D team achieves higher performance across all demographic groups except for white males and females (98.40% vs. 97.90% AVG, and 97.60% vs. 96.50% AVG, respectively). We believe this is produced due to the class imbalance in the real data for this demographic group, which also leads to achieving the best performance for the case of training only with real data, due to overfitting. Similarly, the ADMIS team achieves higher performance across all demographic groups except for black males (98.00% vs. 97.10% AVG). Moreover, when we analyze the scenario of training only with synthetic data, not only does the overall performance increase, but the bias between demographic groups is also reduced (0.82% vs 4.40% STD for Sub-Tasks 1.1 and 1.2, and 0.75% vs 1.81% STD for Sub-Task 1.3), which implies a better model in terms of fairness across all demographic groups. This is mainly produced as with synthetic data we can control how balanced the data is with respect to these groups. Finally, it is noteworthy to compare the best results achieved in Sub-Task 1.2, i.e., unconstrained synthetic data, and Sub-Task 1.3, i.e., constrained synthetic + real data. The ID R&D team achieves 96.73% TO in Sub-Task 1.2, whereas ADMIS achieves 96.50% TO in Sub-Task 1.3, showing that unlimited synthetic data for training can even outperform FR systems trained with limited synthetic + real data. These results motivate the use of synthetic data for demographic bias mitigation, improving at the same time privacy as no real identities are seen by the network.
Sub-Task 2.1 (Overall Improvement): Synthetic Data (Constrained) | |||||
Pos. | Team | AVG [%] | FNMR@FMR=1% | AUC [%] | GAP [%] |
1 | OPDAI | 91.93 | 17.63 (2) | 97.30 (2) | 3.09 |
2 | ID R&D | 91.86 | 10.36 (1) | 97.48 (1) | 2.99 |
3 | ADMIS | 91.19 | 20.41 (3) | 97.04 (3) | 2.78 |
4 | K-IBS-DS | 91.05 | 24.87 (6) | 96.09 (6) | 2.60 |
5 | CTAI | 90.59 | 21.88 (4) | 96.40 (5) | -1.94 |
6 | BOVIFOCR-UFPR | 89.97 | 24.04 (5) | 96.70 (4) | 3.71 |
Sub-Task 2.2 (Overall Improvement): Synthetic Data (Unconstrained) | |||||
Pos. | Team | AVG [%] | FNMR@FMR=1% | AUC [%] | GAP [%] |
1 | Idiap-SynthDistill | 93.50 | 9.17 (1) | 97.17 (4) | -0.05 |
2 | ADMIS | 92.92 | 14.45 (3) | 97.76 (1) | 0.21 |
3 | OPDAI | 92.04 | 16.48 (4) | 97.41 (3) | 3.00 |
4 | ID R&D | 91.86 | 10.36 (2) | 97.48 (2) | 2.99 |
5 | K-IBS-DS | 91.61 | 22.48 (6) | 96.54 (5) | 1.96 |
6 | CTAI | 90.59 | 21.88 (5) | 96.40 (6) | -1.94 |
Sub-Task 2.3 (Overall Improvement): Synthetic + Real Data (Constrained) | |||||
Pos. | Team | AVG [%] | FNMR@FMR=1% | AUC [%] | GAP [%] |
1 | K-IBS-DS | 95.42 | 9.49 (5) | 98.14 (6) | -2.15 |
2 | OPDAI | 95.23 | 7.54 (1) | 98.70 (1) | -0.52 |
3 | CTAI | 94.56 | 8.85 (4) | 98.41 (3) | -6.01 |
4 | CBSR-Samsung | 94.20 | 8.62 (3) | 98.17 (4) | -4.40 |
5 | ADMIS | 94.15 | 10.99 (6) | 98.46 (2) | -1.10 |
6 | ID R&D | 94.05 | 8.00 (2) | 98.16 (5) | 0.07 |
Sub-Task | Team | Model | AgeDB | BUPT | CFP-FP | ROF | AVG [%] | STD [%] | GAP [%] |
2.1 | OPDAI | Proposed | 94.54 | 93.42 | 92.51 | 87.22 | 91.93 | 2.81 | 3.09 |
Baseline | 96.78 | 93.47 | 97.29 | 91.39 | 94.73 | 2.42 | |||
2.2 | Idiap SynthDistill | Proposed | 96.72 | 93.85 | 96.14 | 87.31 | 93.5 | 3.74 | -0.05 |
Baseline | 97.6 | 97.22 | 95.1 | 84.14 | 93.52 | 5.5 | |||
2.3 | K-IBS-DS | Proposed | 96.89 | 96.88 | 97.51 | 90.39 | 95.42 | 2.91 | -2.16 |
Baseline | 96.2 | 93.76 | 95.97 | 87.58 | 93.38 | 3.48 |
@@ V-B Task 2: Overall Improvement
Table VIII provides the results achieved by participants in Task 2, focusing not only on bias mitigation but also other challenges in FR such as age, pose, and occlusions. Teams are ranked in descending order based on the average verification accuracy across the four databases. Notably, in all sub-tasks, the AVG is lower than the achieved in Task 1 for the BUPT-BalancedFace [52] database, showing the additional challenges introduced by the AgeDB [53], CFP-FP [54], and ROF [55] databases. This trend can also be observed in GAP results, which tend to be worse for Sub-Tasks 2.1 and 2.2 compared to Sub-Tasks 1.1 and 1.2, suggesting that it is far more difficult to emulate the conditions of these real databases with synthetic data by itself. For example, in Sub-Task 2.1, although the top-3 teams achieve high AVG results, they exhibit a considerable positive GAP value (i.e., OPDAI 91.93% AVG, 3.09% GAP; ID R&D 91.86% AVG, 2.99% GAP; ADMIS 91.19% AVG, 2.78% GAP), showing that a FR model trained only with real data (i.e., CASIA-WebFace) adapts better to adverse image conditions such as aging, pose, or occlusions. Focusing on Sub-Task 2.2, the team ranked top-1, i.e., Idiap-SynthDistill, achieves much better results compared to the best result of Sub-Task 2.1 (i.e., 93.50% vs. 91.93% AVG), proving that unlimited synthetic data can further improve the performance of the system. Finally, in Sub-Task 2.3, most teams report better AVG and higher negative GAP values (e.g., K-IBS-DS achieves 95.42% AVG, -2.15% GAP), proving again that synthetic data combined with real data can alleviate existing limitations within FR technology.
Furthermore, Table IX compares the performance across the different evaluation databases, highlighting the variations between the baseline models, which were trained only on real data, and the models proposed by the top-ranked teams for the sub-tasks of Task 2.In Sub-Task 2.1, where the model proposed by OPDAI is trained with limited synthetic data due to competition rules, it underperforms compared to the baseline across all databases (3.09% GAP). This suggests that 500,000 synthetic samples are insufficient to fully replace real data. Moreover, in Sub-Task 2.2, where training synthetic data is unlimited, the performance of the model proposed by Idiap-SynthDistill is almost identical to the baseline (-0.05% GAP). This suggests that a large synthetic dataset can closely replicate the distribution of real data. We can also observe that in AgeDB and BUPT databases, the baseline still outperforms the proposed model, which indicates that synthetic data does not fully capture the differences between aging or demographic groups. Finally, in Sub-Task 2.3, which combines real and unlimited synthetic data, the model proposed by K-IBS-DS outperforms the baseline in all databases (-2.16% GAP). This supports the idea of all our previous experiments, that synthetic data complements real data, improving generalization.